(Bayes’ Theorem) It is known that 4% of the population has the dreaded cooties virus. After being exposed on the playground, you go to the doctor to get tested for cooties. The test indicates that you have the virus. The test for cooties has a miss rate of 2% and a false positive rate of 8%. What is the probability that you actually have the cooties virus?
Answer #1

P(tested positive)=P(have virus and tested positive)+P(not have virus and tested positive)

=0.04*(1-0.02)+0.96*0.08=0.1160

hence probability that you actually have the cooties virus given tested positive)

=P(have virus and tested positive)/P(tested positive)=0.04*(1-0.02)/0.1160=0.337931